OpenStudy (anonymous):

Kurt takes out a $4000 loan to buy a car. His monthly payment is $120. Create a table and equation to represent how much money Kurt owes after n months. can someone help me write the equation i would use with n in it please

OpenStudy (anonymous):

@satellite73

OpenStudy (tkhunny):

You'll need an interest rate. If there is no interest... After 1 month 4000 - 120 After 2 months 4000 - 120 - 120 After 3 months 4000 - 120 - 120 - 120 After 4 months 4000 - 120 - 120 - 120 - 120 Are we seeing a pattern?

OpenStudy (anonymous):

its p(n)+1 all the way to 10 but i dont know the p(n) part

OpenStudy (tkhunny):

Okay, let's try again. You'll need an interest rate. If there is no interest... After 1 month 4000 - 120 = 4000 - 120(1) After 2 months 4000 - 120 - 120 = 4000 - 120(2) After 3 months 4000 - 120 - 120 - 120 = 4000 - 120(3) After 4 months 4000 - 120 - 120 - 120 - 120 = 4000 - 120(4)